Solve the ff. turn them to y form
1/3x=y+2
3x-2y+8=0
4x-10y=0
3/4x+2y=8
2/3y=-x/2-5/3x
1/2x+3/2x=2y+5-3y

Answer by checkley77(12844)

1/3x=y+2
y=x/3-2
----------------------------
3x-2y+8=0
-2y=-3x-8
y=-3x/-2-8/-2
y=3x/2+4
-----------------------------
4x-10y=0
-10y=-4x
y=-4x/-10
y=4x/10
y=2/5
---------------------------
3/4x+2y=8
2y=-3x/4+8
y=-3x/8+8/2
y=-3x/8+4
---------------------------
I'll leave the rest for you to practice with.